OpenStudy (anonymous):

solve \[\sqrt{3x-2}\le2 \]

OpenStudy (anonymous):

@mathstudent55 hey come

OpenStudy (anonymous):

what would the answer be? i gotta zip through these i dont want you to have to leave i got two others after this one and then im done :D

OpenStudy (mathstudent55):

Change the inequality into an equation: \(\sqrt{3x - 2} = 2 \) and solve for x. Start be squaring both sides.

OpenStudy (anonymous):

i dont know what to do can you please just solve it for me

OpenStudy (anonymous):

@mathstudent55 can you come back and solve it for me pleaseee

OpenStudy (anonymous):

please math im really depending on you your like crazy smart <3

OpenStudy (mathstudent55):

3x - 2 = 4 3x = 6 x = 2 There is one point of interst, x = 2. Now look at the root. Solve 3x - 2 = 0 3x = 2 x = 2/3 Here is another point of interest, x = 2/3. Now plot them on a number line.

OpenStudy (mathstudent55):